OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible (x^2 +9)e^3x+8 dx

zepdrix (zepdrix):

\[\large \int\limits (x^2+9)e^{3x+8}\;dx\] So is this what we're dealing with? Your exponent doesn't have brackets, so I just wanted to make sure.

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

See how we have a `polynomial` and an `exponential`? Think about the properties of those two ..things. What happens when you differentiate a polynomial over and over? It will decrease in degree right? The powers will keep shrinking until nothing is left. The exponential however will stay the same, (but maybe we end up with some extra junk due to the chain rule). We want our U to be the polynomial. It will decrease in degree, until it eventually deteriorates, leaving us with an easy integral. \[\large u=x^2+9 \qquad \qquad dv=e^{3x+8}dx\] Can you find the \(\large du\) and \(\large v\) ?

OpenStudy (anonymous):

du=2x right

OpenStudy (anonymous):

how do I find v

zepdrix (zepdrix):

Do you know how to integrate this? \[\large \int\limits e^{3x}dx\]

OpenStudy (anonymous):

no is it 3e^x

OpenStudy (anonymous):

(1/3)*e^(3x);

OpenStudy (anonymous):

is this how I set it up (x^2+9)1/3e^3x -1/3e^3x (2x)

OpenStudy (anonymous):

first part is (x^2+9)*(1/3)*e^(3x+8)

OpenStudy (anonymous):

in the second part you must using integration by parts

OpenStudy (anonymous):

so the second part isnt 1/3e^3x (2x)

OpenStudy (anonymous):

yes, you must integrate x*e^(3x+8) by parts

OpenStudy (anonymous):

x1/3e^3x+1 ?